### v3

# The Adoption and Efficacy of Large Language Models: Evidence From Consumer Complaints in the Financial Industry

The Adoption and Efficacy of Large Language Models:
Evidence From Consumer Complaints in the Financial Industry
Minkyu Shin Minkyu Shin: Assistant Professor of Marketing, City University of Hong Kong, minkshin@cityu.edu.hk. Jin Kim: Postdoctoral Research Associate, Northeastern University, jin.kim1@northeastern.edu. Jiwoong Shin: Professor of Marketing, Yale School of Management, jiwoong.shin@yale.edu. We thank the Department of Marketing at City University of Hong Kong for their financial support, enabling us to fully use the AI detection API. We also thank Peter Arcidiacono, Anthony Dukes, Hortense Fong, Christoph Fuchs, Soheil Ghili, Thomas L. Griffiths, Liang Guo, Vanya Klenovskiy, Yanzhi Li, Grant Packard, Stefano Puntoni, K Sudhir, Kosuke Uetake, Gal Zauberman, Joann Peck for their constructive feedback, and seminar participants at Annual Business & Generative AI Workshop by Wharton, Korea University, KAIST, INFORMS Marketing Science Conference, and Marketing Exchange Forum.    Jin Kim    Jiwoong Shin
(October 31, 2024)
Large Language Models (LLMs) are reshaping consumer decision-making, particularly in communication with firms, yet our understanding of their impact remains limited. This research explores the effect of LLMs on consumer complaints submitted to the Consumer Financial Protection Bureau from 2015 to 2024, documenting the adoption of LLMs for drafting complaints and evaluating the likelihood of obtaining relief from financial firms. Utilizing a leading AI detection tool, we analyzed over 1 million complaints and identified a significant increase in LLM usage following the release of ChatGPT. We establish a causal relationship between LLM usage and an increased likelihood of obtaining relief by employing instrumental variables to address endogeneity in LLM adoption. Experimental data further support this link, demonstrating that LLMs enhance the clarity and persuasiveness of consumer narratives. Our findings suggest that facilitating access to LLMs can help firms better understand consumer concerns and level the playing field among consumers. This underscores the importance of policies promoting technological accessibility, enabling all consumers to effectively voice their concerns.
Keywords: Large language models; Consumer complaint; Generative AI Adoption; Consumer finance;
                    Financial Literacy; The Impact of AI
1 Introduction
Large Language Models (LLMs) and generative AI are profoundly influencing everyday human activities across a broad spectrum, particularly by enhancing the precision and efficacy of language in human interactions. This impact stems from the fundamental nature of LLMs, which are centered around language—a crucial element in human communication. Tools like ChatGPT are revolutionizing persuasive dialogue by aiding users in drafting and refining messages. This assistance enhances the clarity and persuasiveness of communication, which is crucial for effective interaction. For instance, recent studies show that academic researchers utilize LLMs to improve the clarity and impact of their writing, increasing reader engagement and comprehension, particularly when writing skills do not match the depth of their insights Liang et al. (2024). Turning to the business context, these capabilities of LLMs are particularly relevant in consumer-initiated communications, such as complaints, where the clarity and persuasiveness of consumer messaging are critical for resolving disputes and securing favorable outcomes from firms. In light of this, our research examines the role of LLMs, such as ChatGPT, in enhancing consumer-firm communications. We focus on two main aspects: the adoption of LLMs by consumers, particularly whether their usage in communications is increasing, and the efficacy of these models in enhancing the persuasiveness of communications, thereby potentially improving outcomes for consumers.
Evidence of LLM adoption among consumers varies widely according to different surveys. For example, a large-scale survey conducted between November 2023 and January 2024 indicated that ChatGPT is pervasive among various occupations in Denmark, with usage rates ranging from 79% among software developers to 34% for financial advisors Humlum and Vestergaard (2024). In contrast, a survey by the Pew Research Center in July 2023 reported a comparatively lower adoption rate of 18% among Americans, which only marginally increased to 23% by February 2024 Park and Gelles-Watnick (2023); McClain (2024). Notably, most existing studies on LLM adoption rely primarily on survey data Bick et al. (2024). Our research contributes as one of the first empirical studies to document adoption patterns using a dataset of more than 1 million consumer complaints spanning nearly a decade, from March 2015 to March 2024. Moreover, by leveraging ZIP code information in the dataset, we trace the trajectory of these adoption patterns and examine their heterogeneity across various sociodemographic groups.
Furthermore, the effectiveness of consumers using LLMs to obtain better responses from firms has yet to be empirically evaluated. While LLMs can enhance the effectiveness of communications by generating content that is coherent, informative, and formal Wu et al. (2022), they might also diminish effectiveness by producing verbose or redundant content Saito et al. (2023), or by failing to capture the subtle nuances of human emotion, which are essential for effective interpersonal communication Shahriar and Hayawi (2023). Additionally, the effective use of LLMs requires substantial domain expertise. Crafting precise prompts often demands significant experience and a thorough understanding of the specific communication context, while assessing the output also necessitates adequate knowledge of the relevant domain. For example, in the finance industry, substantial evidence of a financial illiteracy problem suggests that many consumers may lack the necessary domain knowledge required for effectively leveraging LLMs Lusardi and Mitchell (2014); Fernandes et al. (2014); Finke et al. (2017). This deficiency indicates that, even if LLMs can enhance the clarity or linguistic sophistication of consumer messages, a lack of domain knowledge may prevent consumers from using them effectively to persuade financial firms and secure favorable outcomes. Therefore, whether and to what extent LLMs impact the actual outcomes in such scenarios remains an open question.
To investigate the adoption and efficacy of using LLMs in consumer communication, we analyzed the consumer complaint database from the Consumer Financial Protection Bureau (CFPB). This dataset serves as an ideal testbed, providing a large-scale, historical, and detailed view of two-sided consumer-firm communication. Our data spans nearly a decade from 2015 to 2024, including the pivotal release of ChatGPT on November 30, 2022. This allows us to analyze how both consumer complaints and firm responses change before and after the introduction of the most popular LLM. The dataset includes over 1.1 million complaints, each containing rich details of the consumer’s issue and the firm’s response, including the relief decisions made by firms. We also leverage ZIP code information for each complaint to merge data from the American Community Survey (ACS) database, enabling us to study the role of sociodemographic factors.
The consumer finance domain, with its inherent communication challenges, is well-suited for examining the impact of LLMs on consumer-firm communications. Financial illiteracy Lusardi and Mitchell (2014), along with the complexity of industry jargon and regulatory language, poses significant challenges for consumers — particularly those with limited English proficiency— in articulating their concerns effectively. This raises the timely question of whether AI tools like LLMs can assist vulnerable consumers in overcoming these challenges Hermann et al. (2023). This scenario, where consumers are motivated by significant personal stakes to enhance their communications, offers an opportunity to evaluate the potential of LLMs to improve effectiveness in consumer-firm communications.
In our analysis of the Consumer Financial Protection Bureau’s extensive dataset of consumer complaints, we document empirical patterns of LLM adoption and examine how these patterns vary across sociodemographic groups. Leveraging these sociodemographic variations as instrumental variables, we provided causal evidence that LLM usage increases the likelihood of obtaining relief. To validate this causal relationship, we conducted controlled laboratory experiments, which confirm that LLM usage to refine and polish consumer complaints significantly increases the probability of receiving a favorable response. These improvements do not require consumers to possess domain-specific expertise or a deep understanding of the recipient’s preferences, as LLMs autonomously optimize key linguistic features Niven and Kao (2019); Manning et al. (2020). This process, as corroborated by our experiments, simultaneously enhances clarity, coherence, fluency, and formality Liu et al. (2019); Chen et al. (2022). This suggests that LLMs can simultaneously enhance multiple features, accommodating the heterogeneous preferences of complaint handlers and increasing the likelihood of favorable responses in line with communication accommodation theory Giles (1979). This comprehensive approach, incorporating both observational and experimental studies, demonstrates the practical benefits of LLMs in enhancing consumer advocacy through clearer and more polished messages.
2 Analysis of LLM Adoption and Efficacy Using the CFPB Dataset
2.1 Datasets & AI Detection Tool
CFPB dataset
The CFPB complaint dataset provides granular information on consumer grievances related to a wide spectrum of financial products and services. A key feature of this dataset is the inclusion of detailed consumer narratives, where individuals describe their experiences in their own words. Unlike informal and often brief reviews on platforms such as Yelp or Amazon, complaints submitted to the CFPB are processed through a formal government channel, positioning them as official documents with the potential to trigger regulatory action or investigations. Given this regulatory context, consumers are likely to craft their complaints with greater care and precision, as these submissions can carry serious implications for firms. This stands in stark contrast to the opinion-oriented, casual nature of reviews on commercial platforms. Moreover, the CFPB’s regulatory framework mandates that firms respond to complaints, offering consumers a clear path to potential redress, which may include refunds, corrections, or other forms of relief based on the outcome of the resolution process.
Beyond the consumer narratives, the dataset includes a comprehensive set of structured variables that provide additional insights into the nature and resolution of the complaints. Each complaint is systematically categorized by issue type (e.g., “Improper use of your report” or “Incorrect information on your report”) and product type (e.g., “Debt collection” or “Credit reporting”). The dataset also records firm responses, categorized as “Closed with monetary relief,” “Closed with non-monetary relief,” or ”Closed with explanation,” offering a clear view of how firms address consumer grievances. In our analysis, we define a successful complaint outcome as one that results in either monetary or non-monetary relief, consistent with prior literature Dou and Roh (2019); Dou et al. (2023). [^1] This definition is particularly relevant in the financial context where distinctions between monetary and non-monetary relief are nuanced. Non-monetary relief, although not immediately quantifiable as direct financial compensation, can yield substantial economic benefits. For example, offering a foreclosure alternative like loan modification or forbearance can prevent home loss and the associated displacement costs—providing significant economic value to the consumer without direct financial payment. Detailed descriptions of all variables are provided in the Online Appendix, section 1A. Moreover, complaints are geocoded by ZIP code level, allowing for the identification of regional patterns in consumer dissatisfaction across the U.S. The ZIP code data further enables the integration of sociodemographic information from the ACS, facilitating a more granular analysis of how LLM adoption varies across different demographic and socioeconomic groups.
ACS dataset
We used the 2017–2021 ACS 5-year estimates to collect representative sociodemographic information at the ZIP code level in order to examine the heterogeneous adoption of LLMs following the release of ChatGPT in November 2022. For each ZIP code, we included key variables such as employment rate, education level, median income, total number of households, English proficiency, and internet access. These variables were chosen to capture cross-sectional variation in sociodemographic characteristics across different regions. While a detailed explanation of each variable’s calculation is provided in the Online Appendix, section 1B, we essentially followed the Census Bureau’s methodology for generating regional profiles.
AI Detection Tool
[…]
2.2 Consumer Adoption of LLMs for Writing Complaints in Financial Industry
Figure 1 presents a time trend of the monthly proportion of complaints that the AI detection tool identifies as likely written by an LLM. Specifically, complaints with an AI Score $\geq$ 99% are hereafter referred to as “Likely-AI” complaints. [^3] This threshold is determined based on an examination of the AI score distribution for all complaints, which exhibits a bimodal pattern: one peak occurs above 99%, while a much more pronounced peak is observed below 5%. More details are provided in the Online Appendix, section 1D. In the months immediately following the initial release of ChatGPT (November 30, 2022), we observe a sharp increase in the proportion of Likely-AI complaints, from near 0% up to 9.8%.
Figure 1: LLM Adoption Observed in the CFPB dataset
Note. This figure displays the monthly percentage of complaints identified as Likely-AI from a dataset of 1,134,512 CFPB complaints between March 2015 and March 2024. A vertical red dotted line indicates the release of ChatGPT on November 30, 2022. After this release, the proportion of Likely-AI complaints steadily increased, reaching a peak of 9.8% by March 2024.
It is also notable that before the release of ChatGPT, the proportion of Likely-AI complaints was very close to zero. Although there were some instances where the AI detection tool incorrectly labeled a complaint as likely written by LLMs before ChatGPT was available (i.e., false positive), the rate was consistently negligible, near 0%. When applying less stringent AI Score thresholds (e.g., 80% or 70%) to define Likely-AI complaints, the rate of false positives increases. However, the surge in the share of Likely-AI complaints remains significant (see Online Appendix, section 1D). Consequently, we use a conservative threshold of 99% AI Score in our analysis to reliably identify complaints likely generated or assisted by LLMs.
Figure 2: LLM Adoption Patterns Varying Across Regions Based on English Proficiency
Note. This figure displays regional LLM adoption patterns, segmented by English proficiency quintiles derived from ACS data. Each panel corresponds to a different quintile of 5-digit ZIP codes. The top panel represents the 1st quintile, featuring regions with the lowest language barriers (4% of people with limited English proficiency on average), while the bottom panel represents the 5th quintile, with the highest (60% of people with limited English proficiency on average).
Furthermore, we observe notable regional heterogeneity in adoption patterns linked to English proficiency. For this analysis, we divided regions into five quintiles (Q1 to Q5) based on the percentage of people with limited English proficiency, using ACS data. Temporal trends were then examined separately for each quintile, as shown in Figure 1, except that the percentages are calculated based on each regional sample. As illustrated in Figure 2, each panel corresponds to one quintile, allowing us to compare the trends across regions with varying English proficiency levels. Notably, adoption was initially concentrated in regions with greater language barriers (4th and 5th quintiles), becoming more uniformly distributed across all regions by 2024, indicating that factors beyond language barriers, such as increased awareness, likely spurred this expanded adoption. These findings suggest that consumers from diverse sociodemographic backgrounds might have adopted LLMs differently when submitting complaints to financial firms. This cross-sectional variation in adoption rates served as one of the main instrumental variables in our analysis of LLM efficacy.
2.3 The Impact of Using LLMs on Getting Relief from Financial Firms
Having identified the Likely-AI complaints, we tested whether firms would respond more positively to those complaints. We categorized complaints submitted after the release of ChatGPT into two groups: Likely-AI complaints (AI scores $\geq$ 99%) and Likely-Human complaints (AI scores $\leq$ 5%). We then compared the outcome indicating whether the complaint received any form of relief — monetary or non-monetary — from firms between these groups. This comparison reveals a significant difference in the probability of receiving relief (Likely-AI: 49.3% vs. Likely-Human: 39.9%, p < .001). [^4] In comparison, the overall relief probability during the year prior to the release of ChatGPT (from December 2021 to November 2022) was 32.5%. Even after accounting for over 200 categories of complaints (where consumers select specific issues of their complaints), the positive association remains statistically significant.
While these results provide suggestive evidence of a positive association between LLM usage and the likelihood of obtaining relief, it is crucial to address potential biases that may arise from non-random selection into LLM usage. Specifically, consumers who choose to use LLMs may differ systematically from those who do not, potentially possessing unobserved characteristics that could influence firms’ responses to complaints. Such characteristics, often correlated with socioeconomic status, could bias the estimated treatment effect due to endogeneity, implying that the positive association may be less about the usage of LLMs and more about the socioeconomic status of the consumers.
Instrumental variables (IVs) estimation
To address concerns regarding endogeneity and non-random selection into LLM usage, we used two demographic variables from the American Community Survey as our instrumental variables: Internet Access and English Proficiency, both measured at the ZIP code level. [^5] These two instrumental variables are also recommended by the innovative AI-assisted search method for identifying valid instrumental variables, as proposed by Han (2024). These variables are based on 5-year estimates (2017-2021), preceding the widespread release of LLMs such as ChatGPT. The timing of these data is crucial for ensuring the exogeneity of the IVs, as it mitigates concerns that they could be contemporaneously correlated with firms’ decisions to provide relief. Moreover, these IVs are likely to affect only the usage of LLM, without directly influencing relief outcomes. This exclusion restriction may hold particularly when controlling for other demographic variables, such as income. [^6] We control for a range of observable demographic characteristics such as income, education, and employment status that might otherwise confound the relationship between LLM usage and relief outcomes. By conditioning on these covariates, we aim to ensure that the IVs are not simply capturing latent socioeconomic factors that could influence both LLM adoption and firm behavior.
Our two IVs, Internet Access and English Proficiency, influence LLM adoption in opposite directions, thereby mitigating the potential impact of unobserved confounding variables. Internet Access facilitates the technical feasibility of adopting LLMs, correlating positively with higher socioeconomic factors that might independently enhance the likelihood of obtaining relief due to greater resources or knowledge. In contrast, lower English Proficiency necessitates the usage of LLMs as a tool to overcome language barriers, allowing users to compose more effective complaints. This is particularly critical in areas where socioeconomic advantages may not be as prevalent. By using these contrasting IVs, our approach effectively isolates the impact of LLM usage from potential socioeconomic confounders that might otherwise confound the observed relationships, ensuring our findings accurately reflect the direct effects of LLMs on the efficacy of consumer complaints, not just underlying socioeconomic disparities. [^7] See Mogstad et al. (2021), for a discussion on the advantages of using multiple IVs.
To test the impact of our IVs on LLM usage, we estimate the following first-stage regression:
$\vspace{-0.2cm}\ln\left(\frac{P(LLM_{i}=1)}{P(LLM_{i}=0)}\right)=\alpha_{1}Internet_{j_{i}}+\alpha_{2}English_{j_{i}}+\mathbf{X}_{j_{i}}^{\prime}\boldsymbol{\gamma}+\delta_{t_{i}}+\delta_{s_{i}}+\delta_{p_{i}}+\delta_{k_{i}}.$ (1)
For complaints submitted after the release of ChatGPT, this logistic regression models the log-odds that a given complaint $i$ was generated using an LLM, using Internet Access and English Proficiency as the primary independent variables, measured at the ZIP code level ($j_{i}$) from 2017–2021. The model includes controls for sociodemographic factors, $\mathbf{X}_{j_{i}}$, such as income, education, employment, and also accounts for several fixed effects: $\delta_{t_{i}}$ for time to capture temporal trends in LLM adoption; $\delta_{s_{i}}$ for state-level effects; $\delta_{p_{i}}$ for financial product categories (e.g., credit cards, mortgages); and $\delta_{k_{i}}$ for complaint issue types (e.g., billing, fraud). These fixed effects help isolate the exogenous influence of the IVs on LLM usage by controlling for heterogeneity across multiple dimensions.
Table 1: First-Stage Regression Results
Dependent variable: Likely-AI Dummy
(1) (2) (3)
Proportion of Households Without Internet Access --0.649∗∗    --0.884∗∗∗
(0.243) (0.248)
Proportion of People With Less English Proficiency      0.249∗∗∗      0.297∗∗∗
(0.063) (0.064)
Year-Month Fixed Effect Y Y Y
State Fixed Effect Y Y Y
Complaint Issue Fixed Effect Y Y Y
Finance Product Fixed Effect Y Y Y
Demographic variables Included Y Y Y
Observations 244,407 244,407 244,407
Unit of Observations Each Complaint ($i$)
Sample Period Dec-2022 $\sim$ Mar-2024
Note. ∗∗p $<$ .01; ∗∗∗p $<$ .001
Table 1 presents the estimates from our first-stage regressions. Column (1) shows that in ZIP codes with a higher proportion of households without internet access, consumers are less likely to submit complaints classified as Likely-AI. Column (2) suggests that in ZIP codes with greater language barriers, indicated by less English proficiency, consumers are more likely to submit Likely-AI complaints. Column (3) confirms the robustness of the effect of these IVs when both are included simultaneously. The first-stage results demonstrate that our IVs significantly explain the variation in whether consumers submit Likely-AI complaints, and these findings hold even when controlling for various fixed effects and additional covariates.
Leveraging exogenous variation from two IVs in the first stage, we employ the two-stage residual inclusion (2SRI) approach outlined by Terza et al. (2008), Wooldridge (2015), and Song et al. (2020). In this 2SRI approach, which is mainly developed and employed to account for a binary outcome variable, residuals from the first-stage regression ($\hat{v}_{i}$) are used as a control function in the second-stage analysis, as shown in Equation (2). To account for the estimation uncertainty from the first-stage regression, we employ a bootstrap procedure with 1,000 replications to estimate standard errors.
$\ln\left(\frac{P(Relief_{i}=1)}{P(Relief_{i}=0)}\right)=\beta_{1}LLM_{i}+\beta_{2}\hat{v}_{i}+\mathbf{X}_{j_{i}}^{\prime}\boldsymbol{\gamma}+\delta_{t_{i}}+\delta_{s_{i}}+\delta_{j_{i}}+\delta_{p_{i}}$ (2)
The main findings, presented in Table 2, shows that consumers who used LLMs to write their complaints are more likely to receive relief from financial firms, suggesting a causal relationship. Each column in Table 2 represents the results with different residuals from various IVs used in the first stage. [^8] Our findings on the causal effect are robust to different specifications of the two-stage estimation. They remain robust when the first-stage regression is specified as a linear probability model. In addition, they are also robust to the use of generalized residuals Wooldridge (2015); Basu et al. (2018).
Table 2: Second-Stage Regression Results (with the First-Stage Control Function)
Dependent variable: Relief Dummy
(1) (2) (3)
Likely-AI Dummy     1.190∗∗∗     1.318∗∗∗      1.315∗∗∗x
(0.181) (0.182) (0.184)
First-Stage IV Internet Access English Proficiency Both
Year-Month Fixed Effect Y Y Y
State Fixed Effect Y Y Y
Complaint Issue Fixed Effect Y Y Y
Finance Product Fixed Effect Y Y Y
Demographic variables Included Y Y Y
Observations 244,407 244,407 244,407
Unit of Observations Each Complaint
Sample Period Dec-2022 $\sim$ Mar-2024
Note. ∗∗p $<$ .01; ∗∗∗p $<$ .001
To further substantiate our findings, we compare complaints submitted both before and after the release of ChatGPT, focusing on regions where consumers have adopted LLMs versus those where they have not. This additional empirical evidence confirms that regions where consumers have embraced LLMs experience a higher likelihood of obtaining relief compared to those where LLMs have not been widely adopted. This robust result supports the main conclusion of our study and is detailed in the Online Appendix, section 1E.
3 Controlled Lab Experiments: Presentation Effect of LLM
We designed our experiments to assess the impact of LLMs on the effectiveness of consumer complaints, focusing on whether the observed improvements in our CFPB data analysis stem solely from enhancements in presentation style. While LLMs typically enhance presentation, there is a possibility that interactions with these models could have inadvertently or deliberately altered the content itself, potentially impacting outcomes. Our three controlled experiments, involving 911 participants in total, directly address this by isolating the effects of presentation enhancements from potential content changes, which are not easily discernible in observational data.
Participants evaluated consumer complaints that were either systematically edited by an LLM (i.e., ChatGPT, specifically GPT-4) to improve clarity, coherence, or professionalism, or left unedited as controls. To ensure that improvements were strictly stylistic, we selected 20 diverse consumer complaints from the CFPB dataset and kept the complaints’ content to be identical across the edited and unedited versions (see Online Appendix, section 2A). These complaints, modified in one of the specified dimensions based on a pilot study’s findings (see Online Appendix, section 2B), represented a broad spectrum of financial issues. [^9] All complaints had AI scores below 1%, were submitted before ChatGPT’s release, and showed minimal prior AI involvement. This setup yielded 80 complaints: 20 unedited (Control) and 60 edited (20 for each of Coherent, Clear, and Professional). The primary outcome measured was the likelihood that participants would award monetary compensation, allowing for a direct assessment of the impact of LLM presentation enhancements on complaint effectiveness.
In the first experiment (N = 210), we presented each participant with five distinct complaints, randomly selected from the set of 80, ensuring that contents of complaints were never repeated for a given participant (for more details, see Online Appendix, section 2C). Participants rated the likelihood of providing monetary compensation in response to each complaint on a scale from 1 (Not likely at all) to 7 (Extremely likely), and this response served as the main dependent measure of the experiment. Additionally, participants assessed each complaint’s clarity, coherence, and professionalism immediately after making their compensation decision on the following items: “How clear [coherent / professional] is the complaint above? (1 = Not clear [coherent / professional] at all, 7 = Extremely clear [coherent / professional])”. Participants then completed exploratory and demographic measures to complete the study.
Results and Discussion.
Consistent with our observational data, the complaints modified by ChatGPT (Coherent, Clear, or Professional) were more likely to receive hypothetical monetary compensation (M = 3.52, SD = 1.91) than the unedited (Control) complaints (M = 2.78, SD = 1.78), Welch’s t(517) = 5.78, p < .001, as illustrated in Figure 3 (Panel A) and in line with the results from the same comparison in the preregistered linear fixed effects model (b = 0.74, SE = 0.093, t = 7.96, p < .001). These results corroborate the findings from Section 2.3, confirming that LLM-enhanced complaints significantly increase the likelihood of obtaining relief.
Moreover, the results also revealed that complaints edited to improve on specific linguistic feature (i.e., Clear, Coherent, and Professional complaints) were perceived as generally improved in all other linguistic features as well (see Figure 3, Panel B). For example, complaints edited to be clearer exceeded Control (unedited) complaints not only in clarity (b = 1.53, SE = 0.10, t = 14.94, p < .001), but also in coherence (b = 1.61, SE = 0.10, t = 15.60, p < .001) and professionalism (b = 2.23, SE = 0.11, t = 20.48, p < .001). These results further support the notion that LLM editing improves the overall presentation quality of complaints across multiple linguistic features simultaneously (Niven and Kao, 2019; Manning et al., 2020; Liu et al., 2019). This improvement in presentation quality underscores that enhancements in presentation alone—without substantive content changes—significantly influence the effectiveness of complaints. This finding reinforces the idea that the effectiveness of LLMs observed in our CFPB data might predominantly stem from improvements in presentation.
Figure 3: Experimental Results on the Comparison Between Human Complaints and AI Complaints
Note. Panel A shows the mean compensation rating (and its 95% CI). Panel B shows the simultaneous improvement in all three linguistic features (relative to the unedited, Control complaints) for all three types of edited complaints.
Robustness and Relevance Tests
To verify the robustness of our findings, we conducted two additional experiments. The first of these, with a larger sample size (N = 301), focused solely on compensation likelihood by removing the evaluation of linguistic features, which minimized potential bias. This test reaffirmed our initial results, with edited complaints consistently showing a higher probability of receiving compensation. Detailed results can be found in the Online Appendix, 2D.
The second additional experiment (N = 400) was conducted with participants who had prior work experience in the finance industry, thereby enhancing the external validity of our findings. Participants were asked to evaluate complaints based on a specific presentation feature, either clarity or professionalism (see Online Appendix, 2E). The results replicated those of the previous experiments, confirming that the observed effects are robust in more ecologically valid samples and are consistent even when participants focus on different presentation features. The results from all three experiments suggest that LLMs can enhance the effectiveness of complaints primarily by improving overall presentation quality, thereby catering to the diverse preferences of complaint handlers across different linguistic dimensions.
4 Conclusion and Implications
Large language models are among the most transformative advancements in AI technology. Despite their widespread influence, the degree to which consumers adopt LLMs for interacting with firms, as well as the actual efficacy of LLMs in consumer-firm communication, both remain significant empirical questions. This paper contributes as one of the first empirical studies to explore and document the rapid adoption of LLMs by consumers for drafting complaints to financial firms following the release of ChatGPT. Our analysis of over 1 million complaints from the CFPB database shows a significant and sharp increase in LLM usage. To address concerns regarding endogeneity and non-random selection into LLM usage, we employed instrumental variables, Internet Access and English Proficiency, that influence only the usage of LLMs, without directly affecting the probability of obtaining relief. This analysis using IVs helps us establish a causal relationship between Likely-AI complaints and an increased likelihood of obtaining financial relief. Our research also complements the extensive analysis of CFPB data with experimental validation. Through experiments, we demonstrate that the enhancements in presentation quality facilitated by LLMs are crucial in increasing the effectiveness of consumer communications and the likelihood of achieving financial relief.
The findings from this study have substantial implications for consumer advocacy, firm practices, and regulatory policies. By facilitating access to LLM technologies such as integrating LLM-aided tools on their websites where customers submit complaints, firms can enhance consumer satisfaction, particularly aiding those who may struggle with articulating their complaints effectively. Leveraging LLMs to improve communication allows firms to focus on the actual content of disputes rather than being influenced by presentation style. This approach could uncover previously underexplored consumer segments whose concerns were less vocal or inadequately expressed, leading to more effective customer service and complaint resolution strategies. Ultimately, this fosters stronger relationships between firms and consumers, ensuring that genuine issues are addressed more efficiently.
From a policy perspective, equitable access to LLM technology is crucial to prevent exacerbating existing disparities in consumer outcomes. As our research suggests, LLM usage can significantly enhance the likelihood of obtaining relief from firms, primarily due to the improved presentation style, and clarity provided by these tools. However, without equal access, less tech-savvy or financially constrained consumers may find themselves at a disadvantage. Policymakers should consider strategies to broaden access to LLM tools, such as making them available through public libraries, consumer advocacy groups, or other community resources.
Moreover, regulatory bodies like the CFPB should play a proactive role in ensuring that LLM technologies are accessible to all consumers, particularly those from vulnerable groups. By promoting the integration of these tools across socio-economic divides, regulators can help level the playing field and ensure that all consumers have equal opportunities to effectively voice their concerns, enhancing the overall fairness and effectiveness of financial dispute resolutions.